OpenStudy (anonymous):

x^2+5=2x i am trying to figure out this equation and dont know how

OpenStudy (anonymous):

put the 2x to the left side x^2-2x+5=0 now solve for x using the quadratic formula

OpenStudy (anonymous):

i got that far im just confused on the quadratic formula.

OpenStudy (anonymous):

ax^2+bx+c=0 \[-b \pm \sqrt{b^2-4ac}\] and divide that by 2a

OpenStudy (anonymous):

so a would be 1 b 2 c=5

OpenStudy (anonymous):

a=1 b=-2 c=5

OpenStudy (anonymous):

thank you i got it

OpenStudy (anonymous):

no problem